# The defect of a cubic threefold

### 08 Dec 2023 v1

@@ 5.2. Containing a plane or a cubic scroll
We first show that a cubic threefold containing a plane or a rational normal cubic scroll does not have a very good line.
@@
Suppose that $X$ contains a plane or a rational normal cubic scroll. Then for every good line $l\in X$, the corresponding curve $\widetilde{D}_{l}$ is reducible.
@@
Since $X$ contains a plane or a cubic scroll, the Fano variety of lines is reducible (see Remark 5.7). Let $F(X):=\cup F_{i}$, where each $F_{i}$ is an irreducible surface. Recall that $F(X)$ is singular at $[m]\in F(X)$ if and only if the line $m\subset X$ passes through a singular point ([AK77, Cor 1.11] and[BVdV79, Cor 4]). First we will show that $\widetilde{D}_{l}$ is not contained in $\cap F_{i}\subset\mathrm{Sing}(F(X))$. Recall that by definition,
@@
Assume now that $\widetilde{D}_{l}$ is irreducible. Then $\widetilde{D}_{l}$ is contained in a unique component of $F(X),$ say $F_{1}$. For $i\neq 1$, let
@@
Now assume that $Z_{i}$ is a surface. Let $p:\mathbb{L}\rightarrow F(X)$ be the universal line over $F(X)$, with projection $q:\mathbb{L}\rightarrow X$. Note that $p$ is a $\mathbb{P}^{1}$ bundle over $F(X)$, and $q$ is generically finite. Let $p_{i}:\mathbb{L}_{i}\rightarrow F_{i}$, $q_{i}:\mathbb{L}_{i}\rightarrow X_{i}$ be the restrictions to the component $F_{i}$. Then $\mathbb{L}_{i}$ is a $\mathbb{P}^{1}$ bundle over the surface $F_{i}$, and $q_{i}$ is a dominant map to the surface $Z_{i}$. In particular, the generic fiber of $q_{i}$ is $1$-dimensional. Since $Z_{i}\subset X\subset\mathbb{P}^{4}$, the intersection of the good line $l$ and $Z_{i}$ is non empty; pick $z\in Z_{i}\cap l$, and consider the fiber $q_{i}^{-1}(z).$ Since the dimension of the fibers of $q_{i}$ is an upper-semi continuous function, $\dim q_{i}^{-1}(z)\geq 1.$ In particular, there is a $1-$parameter family of lines contained in $F_{i}$ that pass through $z$, parametrised by a curve $C_{z}\subset F_{i}$. Clearly, $C_{z}$ is a component of $\widetilde{D}_{l}$, and $\widetilde{D}_{l}$ is reducible. ∎
Using the equivalence of Theorem 1.2, this immediately gives the following corollary.
Corollary 5.10.
Let $X$ be a cubic fourfold with $\sigma(X)>0$. Then $X$ does not contain a very good line.
@@ 5.3. Assuming non-existence of a very good line
To complete the proof of Theorem 5.1, it remains to prove the converse: if $X$ does not have a very good line, then $\sigma(X)>0$. By [CML09, §3.2], $X$ has a very good line if $\mu(X)\leq 5$, or if the singularities of $X$ are allowable (see also [LSV17, §2]). Note that such cubic threefolds have $\sigma(X)=0$ by Proposition 3.2.
Definition 5.11.
A cubic threefold $X$ has allowable singularities if it has singularities of type $A_{k}$ with $k\leq 5$ or $D_{4}$ singularities.
We thus make the following assumption for the rest of this section:
Assumption 5.12.
The cubic threefold $X$ satisfies $\mu(X)\geq 6,$ there is at least one non-allowable singularity $q\in X$ and $X$ does not have a singularity with corank $3$.
Suppose that $l\subset X$ is a good line, but is not very good. Then by definition the associated curve $\widetilde{D}_{l}\subset F(X)$ has multiple components, but the discriminant curve $D_{l}:=\widetilde{D}_{l}/\tau\subset\mathbb{P}^{2}$ can be either reducible or irreducible. We consider each case separately.
@@
Theorem 5.1 follows from Theorem 5.13 and the following theorem:

### 08 Jan 2024 v2

@@ 5.2. Containing a plane or a cubic scroll
In this subsection, we will show that a cubic threefold containing a plane or a rational normal cubic scroll (and thus $\sigma(X)>0$) does not contain a very good line. This proves one direction of Theorem 5.1.
@@
Suppose that $X$ contains a plane or a rational normal cubic scroll. Then for every good line $l\in X$, the corresponding curve $\widetilde{D}_{l}$ is reducible. In particular, $X$ does not contain a very good line.
@@
Since $X$ contains a plane or a cubic scroll, the Fano variety of lines is reducible (see Remark 5.7). Let $F(X):=\cup F_{i}$, where each $F_{i}$ is an irreducible surface. Recall that $F(X)$ is singular at $[m]\in F(X)$ if and only if the line $m\subset X$ passes through a singular point ([AK77, Cor 1.11] and[BVdV79, Cor 4]). First we will show that $\widetilde{D}_{l}$ is not contained in $\mathrm{Sing}(F(X))$. Recall that by definition,
@@
Assume now that $\widetilde{D}_{l}$ is irreducible. Since $\widetilde{D}_{l}\not\subset\mathrm{Sing}(X)$, it follows that $\widetilde{D}_{l}$ is contained in a unique component of $F(X),$ say $F_{1}$. For $i\neq 1$, let
@@
Now assume that $Z_{i}\subset X$ is a surface. Let $p:\mathbb{L}\rightarrow F(X)$ be the universal line over $F(X)$, with projection $q:\mathbb{L}\rightarrow X$. Note that $p$ is a $\mathbb{P}^{1}$ bundle over $F(X)$, and $q$ is generically finite. Let $p_{i}:\mathbb{L}_{i}\rightarrow F_{i}$, $q_{i}:\mathbb{L}_{i}\rightarrow Z_{i}$ be the restrictions to the component $F_{i}$. Then $\mathbb{L}_{i}$ is a $\mathbb{P}^{1}$ bundle over the surface $F_{i}$, and $q_{i}$ is a dominant map to the surface $Z_{i}$. In particular, the generic fiber of $q_{i}$ is $1$-dimensional. Since $Z_{i}\subset X\subset\mathbb{P}^{4}$, the intersection of the good line $l$ and $Z_{i}$ is non empty; pick $z\in Z_{i}\cap l$, and consider the fiber $q_{i}^{-1}(z).$ Since the dimension of the fibers of $q_{i}$ is an upper-semi continuous function, $\dim q_{i}^{-1}(z)\geq 1.$ In particular, there is a $1-$parameter family of lines contained in $F_{i}$ that pass through $z$, parametrised by a curve $C_{z}\subset F_{i}$. Since $z\in l$, the curve $C_{z}\subset F_{i}$ is a component of $\widetilde{D}_{l}$, and $\widetilde{D}_{l}$ is reducible. ∎
@@ 5.3. Assuming non-existence of a very good line
In order to complete the proof of Theorem 5.1, it remains to prove the converse: if $X$ does not have a very good line, then $\sigma(X)>0$. Note that if $X$ does not have a good line, then $\sigma(X)>0$ by Lemma 5.8, or $X$ has a corank 3 singularity: we thus assume the existence of a good line in what follows. Further, a cubic threefold has a very good line as long as it appears as a hyperplane section of a general cubic fourfold [LSV17, Proposition 2.10]. In particular, any cubic threefold with $\mu(X)\leq 5$ has a very good line. Thus we can assume that $X$ has $\mu(X)\geq 6$.
Suppose that $l\subset X$ is a good line, but is not very good. Then by definition the associated curve $\widetilde{D}_{l}\subset F(X)$ has multiple components, but the discriminant curve $D_{l}:=\widetilde{D}_{l}/\tau\subset\mathbb{P}^{2}$ can be either reducible or irreducible. Theorem 5.1 will then follow from Theorem 5.10, which considers the case where $D_{l}$ is reducible, and Theorem 5.13, which considers the case where $D_{l}$ is irreducible.
@@
Finally, we finish the proof of Theorem 5.1 by considering the case when the quintic plane curve $D_{l}$ is irreducible.